# On the continuous properties for the 3D incompressible rotating Euler equations

### v1

@@
Abstract: In this paper, we consider the Cauchy problem for the 3D Euler equations with the Coriolis force in the whole space. We first establish the local-in-time existence and uniqueness of solution to this system in $B^{s}_{p,r}(\mathbb{R}^{3})$. Then we prove that the Cauchy problem is ill-posed in two different sense: (1) the solution of this system is not uniformly continuous dependence on the initial data in the same Besov spaces, which extends the recent work of Himonas-Misiołek [19, Comm. Math. Phys., 296, 2010] to the general Besov spaces framework; (2) the solution of this system cannot be Hölder continuous in time variable in the same Besov spaces. In particular, the solution of the system is discontinuous in the weaker Besov spaces at time zero. To the best of our knowledge, our work is the first one addressing the issue on the failure of Hölder continuous in time of solution to the classical Euler equations with(out) the Coriolis force.
@@ 1 Introduction
We say that the Cauchy problem
$\begin{cases}\partial_{t}f=F(f),\\ f(0,x)=f_{0}(x).\end{cases}$ (1.1)
@@
    (Continuous Dependence)  The data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ is continuous in the following sense: for any $T_{1}<T$ and $\varepsilon>0$, there exists $\delta=\delta(\varepsilon,u_{0},\widetilde{u}_{0})>0$, such that if $\|u_{0}-\widetilde{u}_{0}\|_{X}\leq\delta$, then $\mathbf{S}_{t}(\widetilde{u}_{0})$ exists up to $T_{1}$ and
    $\|\mathbf{S}_{t}(u_{0})-\mathbf{S}_{t}(\widetilde{u}_{0})\|_{\mathcal{C}([0,T),X)}\leq\varepsilon.$
Furthermore, we say that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ is uniform continuous if for any $T_{1}<T$ and $\varepsilon>0$, there exists $\delta=\delta(\varepsilon)>0$ independent of $u_{0}$ or $\widetilde{u}_{0}$ such that, if $\|u_{0}-\widetilde{u}_{0}\|_{X}\leq\delta$, then $\mathbf{S}_{t}(\widetilde{u}_{0})$ exists up to $T_{1}$ and
$\|\mathbf{S}_{t}(u_{0})-\mathbf{S}_{t}(\widetilde{u}_{0})\|_{\mathcal{C}([0,T),X)}\leq\varepsilon.$
We would like to emphasize that, in this paper, we say that the Cauchy problem (1.1) is well-posed in a Banach space $X$ if it enjoys existence, uniqueness and uniform continuous dependence on the initial data. In particular, we are interested in the lack of continuity with respect to time variable or uniform continuity of solution with respect to the initial data.
@@ 1.1 Known Well/Ill-posedness results
$\begin{cases}\partial_{t}u+\Omega e_{3}\times u+u\cdot\nabla u+\nabla P=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ \mathrm{div\,}u=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ u(t=0,x)=u_{0}(x),&\quad x\in\mathbb{R}^{3}.\end{cases}$ (1.2)
where the vector field $u(t,x):[0,\infty)\times{\mathbb{R}}^{3}\to{\mathbb{R}}^{3}$ stands for the velocity of the fluid, the quantity $P(t,x):[0,\infty)\times{\mathbb{R}}^{3}\to{\mathbb{R}}$ denotes the scalar pressure, and $\mathrm{div\,}u=0$ means that the fluid is incompressible. The constant $\Omega\in\mathbb{R}$ represents the speed of rotation around the vertical unit vector $e_{3}=(0,0,1)$, which is called the Coriolis parameter. Problems concerning large-scale atmospheric and oceanic flows are known to be dominated by rotational effects. By this reason, almost all of the models of oceanography and meteorology dealing with large-scale phenomena include a Coriolis force. For example, an oceanic circulation featuring a hurricane is caused by the large rotation. There is no doubt that other physical effects are of similar significance like salinity, natural boundary conditions and so on. The dispersive effect of rotation in fluid flows has been studied in the literature from various perspectives, see, e.g., geophysical flows [16, 27, 31], life span and asymptotic behaviour in the case of fast rotations [10, 11, 15, 26] and almost global stability [17].
@@
The classical Euler equations.  In the case of $\Omega=0$, (1.2) reduces to the original Euler equations. Kato [21] proved the local well-posedness of classical solution to Euler equations in the Sobolev space $H^{s}(\mathbb{R}^{3})$ for all $s>5/2$. Kato-Ponce [23] extended this result to the Sobolev spaces $W^{s,p}(\mathbb{R}^{3})$ of the fractional order for $s>3/p+1,1<p<\infty$. Chae [7, 8, 9] and Chen-Miao-Zhang [12] gave further extensions to the Triebel-Lizorkin spaces $F_{p,q}^{s}(\mathbb{R}^{3})$ with $s>3/p+1,1<p,q<\infty$ and the Besov spaces $B_{p,q}^{s}(\mathbb{R}^{3})$ with $s>3/p+1$, $1<p<\infty,1\leq q\leq\infty$ or $s=3/p+1,1<p<\infty,q=1$. However, these two kinds of function spaces are only in the $L^{p}(1<p<\infty)$-framework since the Riesz transform is not bounded on $L^{\infty}$. The currently-known best result on the local existence was given by Pak-Park [30] in the Besov space $B_{\infty,1}^{1}(\mathbb{R}^{3})$. Guo-Li-Yin [18] proved the continuous dependence of the Euler equations in the space $B_{p,q}^{s}(\mathbb{R}^{3})$ with $s>3/p+1$, $1\leq p\leq\infty,1\leq q<\infty$ or $s=3/p+1,1\leq p\leq\infty,q=1$. Later, Cheskidov-Shvydkoy [14] proved that the solution of the Euler equations cannot be continuous as a function of the time variable at $t=0$ in the spaces $B^{s}_{r,\infty}(\mathbb{T}^{d})$ where $s>0$ if $2<r\leq\infty$ and $s>d(2/r-1)$ if $1\leq r\leq 2$. Furthermore, Bourgain-Li [4, 5] proved the strong local ill-posedness of the Euler equations in borderline Besov spaces $B^{d/p+1}_{p,r}$ with $(p,r)\in[1,\infty)\times(1,\infty]$ when $d=2,3$. Subsequently, Misiołek-Yoneda [28] studied the borderline cases and showed that the 2D Euler equations are not locally well-posed in the sense of Hardamard in the $C^{1}$ space and in the Besov space $B^{1}_{\infty,1}$. Recently, Misiołek-Yoneda [29] showed that the solution map for the Euler equations is not even continuous in the space of Hölder continuous functions and thus not locally Hadamard well-posed in $B^{1+s}_{\infty,\infty}$ with any $s\in(0,1)$. Concerning the non-uniform continuity of the data-to-solution map, we would like to mention that the beautiful results of Himonas-Misiołek [19] covered both the torus $\mathbb{T}^{d}$ and the whole spaces $\mathbb{R}^{d}$ cases. More precisely, they proved that the solution map for the Euler equations in bi(tri)-dimension is not uniformly continuous in Sobolev spaces $H^{s}(\mathbb{T}^{d})$ for $s\in\mathbb{R}$ and in $H^{s}(\mathbb{R}^{d})$ for any $s>0$. Bourgain and Li [6] settled the border line case $s=0$.
@@
The Euler equations with the Coriolis force.  In the case of $\Omega\neq 0$, as it is already observed in [10, 11, 15], the Euler-Coriolis system (1.2) exhibits a dispersion phenomenon which is due to the presence of the Coriolis force $\Omega e_{3}\times u$. For large Coriolis parameter $|\Omega|$, Dutrifoy [15] showed the asymptotics of solutions to vortex patches or Yudovich solutions as the Rossby number goes to zero for some particular initial data, and gave a lower bound on the lifespan $T_{\Omega}$ of the solution to (1.2) as $T_{\Omega}\gtrsim\log\log|\Omega|$. For $u_{0}\in H^{s}(\mathbb{R}^{3})$ with $s>5/2$, Koh-Lee-Takada [25] proved that there exists a unique local in time solution $u$ for (1.2) with $\Omega\in\mathbb{R}$ in the class $\mathcal{C}([0,T];H^{s}(\mathbb{R}^{3}))\cap\mathcal{C}^{1}([0,T];H^{s-1}(\mathbb{R}^{3}))$ (see also [33, 34]). Moreover, assuming that $s>7/2$, they showed that their solutions can be extended to long-time intervals $\left[0,T_{\Omega}\right]$ provided that the speed of rotation is large enough. V. Angulo-Castillo and L.C.F Ferreiragave [1] further gave extensions to the critical Besov space $B^{5/2}_{2,1}(\mathbb{R}^{3})$. Jia-Wan [20] proved the long time existence of classical solutions to (1.2) for initial data in the Sobolev space $H^{s}$ with $s>5/2$ by a weaker assumption on the lower bound of $|\Omega|$. However, the above mentioned results for (1.2) only involve the short-time existence, uniqueness for any speed of rotation and long-time existence for high speed of rotation. We should mention that the dispersion relation underlying the linear system
@@
$\begin{cases}\partial_{t}u+\Omega e_{3}\times u+\nabla P=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ \mathrm{div\,}u=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ u(t=0,x)=u_{0}(x),&\quad x\in\mathbb{R}^{3}.\end{cases}$
@@
is given by $\Lambda(\xi)=\Omega\frac{\xi_{3}}{|\xi|}$. Compared with the classical Euler equations, we would like to emphasize that, this dispersive mechanism exhibits $O(t^{-1})$ decay rate in $L^{\infty}$-norm (see [32]) and is strongly anisotropic and degenerate. It is still surprising that, in the absence of viscosity, the Coriolis force alone is sufficient to stabilize the solutions globally in time in the full 3D setting (see [17]). It is found that the strong rotational effect enhances the temporal decay rate of a certain norm of the velocity (see [2]).
@@ 1.2 Motivation
A fundamental challenge in mathematical physics is to understand the behavior of solutions for the complex rotating fluids. From the PDE’s point of view, it is crucial to know if an equation which models a physical phenomenon is well-posed in the Hadamard’s sense: existence, uniqueness, and continuous dependence of the solutions with respect to the initial data. In particular, the lack of continuous dependence would cause incorrect solutions or non meaningful solutions. It is known that, the non-uniform continuity of data-to-solution map suggests that the local well-posedness cannot be established by the contraction mappings principle since this would imply Lipschitz continuity for the solution map. For certain types of evolutionary equations, for example, the Korteweg-de Vries (KdV) and nonlinear Schrödinger equations, the flow map is Lipschitz continuous. Paticularly, the KdV equation is shown to be well-posed in $H^{s}(\mathbb{R})$ with $s>-3/4$, then the solution map is Lipschitz on the same $H^{s}(\mathbb{R})$ (see [24]). However, for the Burgers equation $\partial_{t}u+u\partial_{x}u=0,$ which can be the most simple quasilinear symmetric hyperbolic equation, Kato [22] showed that the dependence on initial data of the Burgers equation in $H^{k}$ with $k\geq 2$ is continuous but not Hölder continuous with any prescribed exponent. This example indicates that the continuity of the solution map for a symmetric hyperbolic system is a delicate issue. Chen-Liu-Zhang [13] proved that the solution map of the b-family equation is Hölder continuous as a map from a bounded set of $H^{s}(\mathbb{R}),s>3/2$ with $H^{r}(\mathbb{R})(0\leq r<s)$ topology, to $C([0,T],H^{r}(\mathbb{R}))$ for some $T>0$. In this paper, for the Cauchy problem (1.2), we shall study the continuity of solution with respect to time variable or uniform continuity of solution with respect to the initial data.
Assume that $u_{0}\in B^{s}_{p,r}$, it is known that there exists a solution $u\in\mathcal{C}([0,T];B^{s}_{p,r})$ for the Euler-Coriolis system (1.2). Naturally, we may wonder whether or not the solution $u$ can belong to $\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with some $\alpha\in(0,1)$. Therefore, we are interested in the following question:
$u_{0}\in B^{s}_{p,r}\;\Rightarrow\;\exists 1\;u\in\mathcal{C}([0,T];B^{s}_{p,r})\;\overset{?}{\Rightarrow}\;\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})\quad\text{with}\;\alpha\in(0,1).$
If the initial data $u_{0}$ have more regularity such that $u_{0}\in B^{s^{\prime}}_{p,r}$ for some $s^{\prime}>s$, by the interpolation argument, we can deduce that $u\in\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with $\alpha=s^{\prime}-s$. In this paper, we will show that there exits initial data $u_{0}\in B^{s}_{p,r}$ such that the corresponding solution of the Euler-Coriolis system (1.2) can not belong to $\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with any $\alpha\in(0,1)$.

### v2

@@
Abstract: In this paper, we consider the Cauchy problem for the 3D Euler equations with the Coriolis force in the whole space. We first establish the local-in-time existence and uniqueness of solution to this system in $B^{s}_{p,r}(\mathbb{R}^{3})$. Then we prove that the Cauchy problem is ill-posed in two different sense: (1) the solution of this system is not uniformly continuous dependence on the initial data in the same Besov spaces, which extends the recent work of Himonas-Misiołek [19, Comm. Math. Phys., 296, 2010] to the more general framework of Besov spaces; (2) the solution of this system cannot be Hölder continuous in time variable in the same Besov spaces. In particular, the solution of the system is discontinuous in the weaker Besov spaces at time zero. To the best of our knowledge, our work is the first one addressing the issue on the failure of Hölder continuous in time of solution to the classical Euler equations with(out) the Coriolis force.
@@ 1 Introduction
$\begin{cases}\partial_{t}u+\Omega e_{3}\times u+u\cdot\nabla u+\nabla P=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ \mathrm{div\,}u=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ u(t=0,x)=u_{0}(x),&\quad x\in\mathbb{R}^{3},\end{cases}$ (1.1)
where the vector field $u(t,x):[0,\infty)\times{\mathbb{R}}^{3}\to{\mathbb{R}}^{3}$ stands for the velocity of the fluid, the quantity $P(t,x):[0,\infty)\times{\mathbb{R}}^{3}\to{\mathbb{R}}$ denotes the scalar pressure, and $\mathrm{div\,}u=0$ means that the fluid is incompressible. The constant $\Omega\in\mathbb{R}$ represents the speed of rotation around the vertical unit vector $e_{3}=(0,0,1)$ and is called the Coriolis parameter. Problems concerning large-scale atmospheric and oceanic flows are known to be dominated by rotational effects. For this reason, almost all of the models of oceanography and meteorology dealing with large-scale phenomena include a Coriolis force. For example, an oceanic circulation featuring a hurricane is caused by the large rotation. There is no doubt that other physical effects are of similar significance like salinity, natural boundary conditions and so on. The dispersive effect of rotation in fluid flows has been studied in the literature from various perspectives, see, e.g., geophysical flows [16, 27, 32], life span and asymptotic behaviour in the case of fast rotations [10, 11, 15, 26] and almost global stability [17].
Before stating related results for (1.1) precisely, we recall the notion of well-posedness in the sense of Hadamard. We say that the Cauchy problem
$\begin{cases}\partial_{t}f=F(f),\\ f(0,x)=u_{0}(x),\end{cases}$ (1.2)
@@
    (Continuous Dependence)  The data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ is continuous in the following sense: for any $u_{0}$, any $T_{1}<T\left(u_{0}\right)$ and any $\varepsilon>0$, there exists $\delta=\delta(\varepsilon,u_{0})$ such that for any $\widetilde{u}_{0}$, if $\left\|u_{0}-\widetilde{u}_{0}\right\|_{X}\leq\delta$ then $\mathbf{S}_{t}(\widetilde{u}_{0})$ exists up to $T_{1}$ and
    $\|\mathbf{S}_{t}(u_{0})-\mathbf{S}_{t}(\widetilde{u}_{0})\|_{\mathcal{C}([0,T_{1}],X)}\leq\varepsilon.$
Furthermore, we say that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ is uniform continuous if for any $T_{1}<T$ and any $\varepsilon>0$, there exists $\delta=\delta(\varepsilon)>0$ independent of $u_{0}$ or $\widetilde{u}_{0}$ such that, if $\|u_{0}-\widetilde{u}_{0}\|_{X}\leq\delta$, then $\mathbf{S}_{t}(\widetilde{u}_{0})$ exists up to $T_{1}$ and
$\|\mathbf{S}_{t}(u_{0})-\mathbf{S}_{t}(\widetilde{u}_{0})\|_{\mathcal{C}([0,T_{1}],X)}\leq\varepsilon.$
We would like to emphasize that, in this paper, we say the Cauchy problem (1.2) is well-posed in a Banach space $X$ if it enjoys existence, uniqueness and continuous dependence on the initial data. We say that the Cauchy problem (1.2) is semi-linearly well-posed (see [30]) for the initial data in $X$ if it is Hadamard well-posed and in addition, the solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ is uniformly continuous from any bounded subset $B\subset X$ into $C\left([0,T];X\right)$. In particular, we are interested in the lack of continuity with respect to time variable or uniform continuity of solution with respect to the initial data.
@@ 1.1 Known Well/Ill-posedness results
The classical Euler equations.  In the case of $\Omega=0$, (1.1) reduces to the original Euler equations. Kato [21] proved the local well-posedness of classical solution to Euler equations in the Sobolev space $H^{s}(\mathbb{R}^{3})$ for all $s>5/2$. Kato-Ponce [23] extended this result to the Sobolev spaces $W^{s,p}(\mathbb{R}^{3})$ of fractional order for $s>3/p+1,1<p<\infty$. Chae [7, 8, 9] and Chen-Miao-Zhang [12] gave further extensions to the Triebel-Lizorkin spaces $F_{p,q}^{s}(\mathbb{R}^{3})$ with $s>3/p+1,1<p,q<\infty$ and the Besov spaces $B_{p,q}^{s}(\mathbb{R}^{3})$ with $s>3/p+1$, $1<p<\infty,1\leq q\leq\infty$ or $s=3/p+1,1<p<\infty,q=1$. However, these two kinds of function spaces are only in the $L^{p}(1<p<\infty)$-framework since the Riesz transform is not bounded on $L^{\infty}$. The currently-known best result on the local existence was given by Pak-Park [31] in the Besov space $B_{\infty,1}^{1}(\mathbb{R}^{3})$. Guo-Li-Yin [18] proved the continuous dependence of the Euler equations in the space $B_{p,q}^{s}(\mathbb{R}^{3})$ with $s>3/p+1$, $1\leq p\leq\infty,1\leq q<\infty$ or $s=3/p+1,1\leq p\leq\infty,q=1$. Later, Cheskidov-Shvydkoy [14] proved that the solution of the Euler equations cannot be continuous as a function of the time variable at $t=0$ in the spaces $B^{s}_{r,\infty}(\mathbb{T}^{d})$ where $s>0$ if $2<r\leq\infty$ and s>d(2/r1)s>d(2/r-1) if $1\leq r\leq 2$. Furthermore, Bourgain-Li [4, 5] proved the strong local ill-posedness of the Euler equations in borderline Besov spaces $B^{d/p+1}_{p,r}$ with $(p,r)\in[1,\infty)\times(1,\infty]$ when $d=2,3$. Subsequently, Misiołek-Yoneda [28] studied the borderline cases and showed that the 2D Euler equations are not locally well-posed in the sense of Hardamard in the $C^{1}$ space and in the Besov space $B^{1}_{\infty,1}$. Recently, Misiołek-Yoneda [29] showed that the solution map for the Euler equations is not even continuous in the space of Hölder continuous functions and thus not locally Hadamard well-posed in $B^{1+s}_{\infty,\infty}$ with any $s\in(0,1)$. Concerning the non-uniform continuity of the data-to-solution map, we would like to mention that the beautiful results of Himonas-Misiołek [19] covered both the torus $\mathbb{T}^{d}$ and the whole spaces $\mathbb{R}^{d}$ cases. More precisely, they proved that the solution map for the Euler equations in bi(tri)-dimension is not uniformly continuous in Sobolev spaces $H^{s}(\mathbb{T}^{d})$ for $s\in\mathbb{R}$ and in $H^{s}(\mathbb{R}^{d})$ for any $s>0$. Bourgain and Li [6] settled the border line case $s=0$.
@@
The Euler equations with the Coriolis force.  In the case of $\Omega\neq 0$, as it is already observed in [10, 11, 15], the Euler-Coriolis system (1.1) exhibits a dispersion phenomenon which is due to the presence of the Coriolis force $\Omega e_{3}\times u$. For large Coriolis parameter $|\Omega|$, Dutrifoy [15] showed the asymptotics of solutions to vortex patches or Yudovich solutions as the Rossby number goes to zero for some particular initial data, and gave a lower bound on the lifespan $T_{\Omega}$ of the solution to (1.1) as $T_{\Omega}\gtrsim\log\log|\Omega|$. For $u_{0}\in H^{s}(\mathbb{R}^{3})$ with $s>5/2$, Koh-Lee-Takada [25] proved that there exists a unique local in time solution $u$ for (1.1) with $\Omega\in\mathbb{R}$ in the class $\mathcal{C}([0,T];H^{s}(\mathbb{R}^{3}))\cap\mathcal{C}^{1}([0,T];H^{s-1}(\mathbb{R}^{3}))$ (see also [34, 35]). Moreover, assuming that $s>7/2$, they showed that their solutions can be extended to long-time intervals $\left[0,T_{\Omega}\right]$ provided that the speed of rotation is large enough. V. Angulo-Castillo and L.C.F Ferreira [1] further gave extensions to the critical Besov space $B^{5/2}_{2,1}(\mathbb{R}^{3})$. Jia-Wan [20] proved the long time existence of classical solutions to (1.1) for initial data in the Sobolev space $H^{s}(s>5/2)$ with a weaker assumption on the lower bound of $|\Omega|$. However, the above mentioned results for (1.1) only involve the short-time existence, uniqueness for any speed of rotation and long-time existence for high speed of rotation. We should mention that the dispersion relation underlying the linear system
@@
$\begin{cases}\partial_{t}u+\Omega e_{3}\times u+\nabla P=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ \mathrm{div\,}u=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ u(t=0,x)=u_{0}(x),&\quad x\in\mathbb{R}^{3},\end{cases}$
@@
is given by $\Lambda(\xi)=\Omega\frac{\xi_{3}}{|\xi|}$. Compared with the classical Euler equations, we would like to emphasize that this dispersive mechanism exhibits $O(t^{-1})$ decay rate in $L^{\infty}$-norm (see [33]) and is strongly anisotropic and degenerate. It is still surprising that, in the absence of viscosity, the Coriolis force alone is sufficient to stabilize the solutions globally in time in the full 3D setting (see [17]). It is found that the strong rotational effect enhances the temporal decay rate of a certain norm of the velocity (see [2]).
@@ 1.2 Motivation
A fundamental challenge in mathematical physics is to understand the behavior of solutions for the complex rotating fluids. From the PDE’s point of view, it is crucial to know if an equation which models a physical phenomenon is well-posed in the Hadamard’s sense: existence, uniqueness, and continuous dependence of the solutions with respect to the initial data. In particular, the lack of continuous dependence would cause incorrect solutions or non meaningful solutions. It is known that, the non-uniform continuity of data-to-solution map suggests that the local well-posedness cannot be established by the contraction mappings principle since this would imply Lipschitz continuity for the solution map. For certain types of evolutionary equations, for example, the Korteweg-de Vries (KdV) and nonlinear Schrödinger equations, the flow map is Lipschitz continuous. Paticularly, the KdV equation is shown to be well-posed in $H^{s}(\mathbb{R})$ with s>3/4s>-3/4, then the solution map is Lipschitz on the same $H^{s}(\mathbb{R})$ (see [24]). However, for the Burgers equation $\partial_{t}u+u\partial_{x}u=0$, which is the most simple quasilinear symmetric hyperbolic equation, Kato [22] showed that the dependence on initial data of the Burgers equation in $H^{k}$ with $k\geq 2$ is continuous but not Hölder continuous with any prescribed exponent. This example indicates that the continuity of the solution map for a symmetric hyperbolic system is a delicate issue. Chen-Liu-Zhang [13] proved that the solution map of the b-family equation is Hölder continuous as a map from a bounded set of $H^{s}(\mathbb{R}),s>3/2$ with $H^{r}(\mathbb{R})(0\leq r<s)$ topology, to $C([0,T],H^{r}(\mathbb{R}))$ for some $T>0$. In this paper, for the Cauchy problem (1.1), we shall study the continuity of solution with respect to time variable or uniform continuity of solution with respect to the initial data.
Assume that $u_{0}\in B^{s}_{p,r}$, based on the local well-posedness theory for the Euler equations (see [3, Theorem 7.1]), we know that there exists a solution $u\in\mathcal{C}([0,T];B^{s}_{p,r})$ for the Euler-Coriolis system (1.1). Naturally, we may wonder whether or not the solution $u$ can belong to $\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with some $\alpha\in(0,1)$. Therefore, we are interested in the following question:
$u_{0}\in B^{s}_{p,r}\;\Rightarrow\;\exists\;u\in\mathcal{C}([0,T];B^{s}_{p,r})\;\overset{?}{\Rightarrow}\;u\in\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})\quad\text{with}\;\alpha\in(0,1).$
If the initial data $u_{0}$ has more regularity, that is $u_{0}\in B^{s^{\prime}}_{p,r}$ for some $s^{\prime}>s$, by the interpolation argument, we can deduce that $u\in\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with $\alpha=s^{\prime}-s$. In this paper, we will show that there exits initial data $u_{0}\in B^{s}_{p,r}$ such that the corresponding solution of the Euler-Coriolis system (1.1) can not belong to $\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with any $\alpha\in(0,1)$.